A business firms sells a good at the price of Rs 450.The firm has decided to reduce the 

price of good to Rs 350.Consequently, the quantity demanded for the good rose from 

25,000 units to 35,000 units. Calculate the price elasticity of demand.


1
Expert's answer
2020-11-10T07:24:22-0500

"E=\frac {Q_2-Q_1}{p_2-p_1}\times \frac{p_1+p_2}{Q_1+Q_2}"


"E=\frac {35,000-25,000}{350-450} \times \frac {450+350}{25,000+35,000}=-1.33"